# Late-Time Evolution of Magnetized Disks in Tidal Disruption Events

### v1

@@ 5 Conclusions
Standard accretion disk models are viscously and thermally unstable in the inner regions of the disk, where radiation pressure dominates over gas pressure (Pringle et al., 1973; Shakura & Sunyaev, 1976; Lightman & Eardley, 1974; Piran, 1978). However, these instabilities are not so far observed in TDEs. A common theoretical approach to stabilize the disk is to modify the viscosity parametrization (Sakimoto & Coroniti, 1981; Taam & Lin, 1984; Mummery & Balbus, 2019), though the most widely used parametrizations are often ad hoc.
In this paper, we examine the $\alpha$-disk model (Shakura & Sunyaev, 1973), under the assumption that magnetic fields are the dominant source of pressure. Given the presence of MRI, we expect the disk to be highly magnetized. Previous research has shown that magnetic fields can indeed stabilize the disk against radiation pressure instabilities (Begelman & Pringle, 2007; Sądowski, 2016; Jiang et al., 2019; Huang et al., 2023; Kaur et al., 2023).
@@
We examined various values for the power law of the linear viscosity and found that, although the surface densities look quite different, the light curves may look very similar when the power law of the linear viscosity is chosen carefully. We found that the time evolution of the UV luminosity for the magnetized viscosity is $L_{\rm m,UV}\propto t^{-5/6}$, and that the light curves for the linear viscosity fortuitously follow the same time evolution when the power law is set to $\mu_{\rm l}=0$.
We fit our model to the well-observed optical and UV light curves of the TDE ASASSN-14li. The evolution of its light curves indicates that the viscous spreading in ASASSN-14li is slow, leading previous studies to propose mechanisms such as angular momentum loss in a magnetized wind (Wen et al., 2023) to limit the disk’s outer spreading. Our results show that the best-fit Shakura-Sunyaev $\alpha$ parameter falls within a theoretically reasonable range, requiring no additional physical effects.
@@
In this paper, we studied accretion disks in a 1D framework, assuming a single pressure component due to magnetic fields. In reality, radiation pressure may also be non-negligible (Jiang et al., 2019) and should be considered alongside magnetic pressure for a more comprehensive model. Furthermore, while we used an approximate estimation for the magnetic field strengths based on an MRI saturation criterion (Pessah & Psaltis, 2005; Begelman & Pringle, 2007) which is consistent with recent radiation MHD simulations (Jiang et al., 2019; Mishra et al., 2022; Huang et al., 2023), alternative saturation criteria (Oda et al., 2009; Begelman & Armitage, 2023) and other MHD instabilities (Das et al., 2018) deserve further investigation and could potentially influence our results. Nevertheless, the magnetized viscosity parametrization adopted in this work remains both simple and more physically motivated compared to other commonly used parametrizations.
@@ Appendix A Separability of Disk Equation
The evolution equation (Eq. 1) with a magnetized $\alpha$-viscosity (Eq. 5) is a nonlinear partial differential equation. Here we show that Eq. 1 can be reduced to two ordinary differential equations: (i) a trivial time-dependent equation and (ii) a nonlinear Emden-Fowler equation. To the best of our knowledge, this has not been demonstrated previously despite the wide applicability of Eq. 1 in high energy astrophysics. There are, however, more specific investigations of separability of Eq. 1 in certain limits, such as self-similar solutions (Rafikov, 2016).
@@
where for the magnetized $\alpha$-viscosity $m=2/7$ and $n=5/7$. We use the following transformation from Pringle (1991): $x\equiv R^{1/2}$, $\tau\equiv 3\nu_{0}t/4$ and $S(\tau,x)\equiv\Sigma R^{3/2}$. Eq. 1 then becomes:
@@
The time-dependent equation is trivial and can be solved easily. For $m=0$, the solution to Eq. 1 has been determined previously (e.g. Metzger et al. 2008), so we focus here on the case $m\neq 0$. In this case, the solution is $T(\tau)=(-m\lambda\tau+C)^{-1/m}$ where $C$ is an arbitrary constant of integration.
The spatial equation is a nonlinear ordinary differential equation, which through a change of variables $w(x)\equiv X^{m+1}x^{2n-3m-2}$ transforms into an Emden-Fowler [^10] The Lane-Emden equation is a well-studied special case of the Emden-Fowler equation. equation (Zaitsev & Polyanin, 2002):
@@
For the magnetized $\alpha$-viscosity we employ, the power of $w$ is $7/9$, and the power of $x$ is $10/9$. Although exact solutions exist for specific power-law values (Zaitsev & Polyanin, 2002), no exact solution has yet been found for these power-law indices.
We have explored other astrophysically interesting power-law viscosity parametrizations. We calculated the values for every combination of gas, radiation, and magnetic pressures ($P\in\left\{P_{\rm gas},P_{\rm rad},P_{\rm mag}\right\}$), with every combination of electron scattering, Kramers, and dust opacities ($\kappa\in\left\{\kappa_{\rm es},\kappa_{\rm Kr},\kappa_{\rm dust}\right\}$). The power-law values of the exact solutions (both continuous and discrete), as well as those of astrophysical interest, are presented in Fig. 9. Unfortunately, no exact solutions currently exist for the values relevant to our study. If in the future such solutions are found, they could provide greater physical intuition for the time evolution of viscously spreading accretion disks.
@@ Appendix B Fitting the Self-Similar Solution
Using the self-similar solution is simpler and more efficient than numerically integrating the disk equations. In §2, we demonstrated that the self-similar solution closely approximates the numerical solution at late times and large radii. In this appendix, we provide a fitting function for the free parameter $\nu_{0,\rm{S}}$ in Eq. 12. For each value of the $\alpha$ parameter, SMBH mass $M_{\bullet}$, and stellar mass $M_{\star}$ satisfying $M_{\bullet}<M_{\rm Hills}$, the self-similar free parameter can be determined using:
@@
Fig. 10 shows the relative error in the light curves at a UV wavelength of $\lambda=250\rm{nm}$, between the numerical and the self-similar solutions:
$\rm{RelativeError(t,\alpha,M_{\bullet},M_{\star})}=100\frac{L_{\nu,\rm{num}}-L_{\nu,\rm{S}}}{L_{\nu,\rm{num}}}\,[\%]$ (B2)
where $t$ is the age of the TDE, $L_{\nu,\rm{num}}$ is the spectral luminosity calculated from the numerical solution, and $L_{\nu,\rm{S}}$ is the spectral luminosity from the self-similar solution. The relative error is larger for SMBH masses near the Hills mass but remains below an order of magnitude in other regions.
@@ Appendix C Comparison To Linear Viscosity with Additional Models
In this appendix, we present additional results. We compare the magnetized $\alpha$-disk with the linear viscosity model, as was done in §3, for two additional scenarios. In the first scenario, we retain the Gaussian initial condition but normalize the linear viscosity prefactor, $\nu_{0,\rm l}$, in the initial conditions, rather than matching it to the magnetized disk after 1 year. Specifically, we set: $\nu_{0,\rm l}=\nu_{\rm mag}\left(r_{\rm c},\Sigma_{0}\right)/r_{\rm c}^{\mu_{\rm l}}$ where $\Sigma_{0}=\Sigma\left(R=r_{\rm c},t=0\right)$ is the initial surface density at the peak of the Gaussian. We compute results for various power-law values $\mu_{\rm l}=\left\{1,0,-3/2\right\}$.
@@
In the second scenario, we add a source function to Eq. 1 instead of using a Gaussian initial condition (while retaining the original normalization $\nu_{0,\rm{l}}$). This approach reflects the slow accumulation of mass from the bound debris into the accretion disk, at a rate given by Eq. 9. For this case, we compute results for $\mu_{\rm l}=\left\{3/2,0,-3/2\right\}$.
@@
In Figs. 11, 12 and 13, we explore these non-fiducial models by plotting the surface density and effective temperature as functions of disk radius, as well as the light curves. For all figures, we use same SMBH, star, and disk parameters as in §3. The best match between the magnetized and linear viscosity models is achieved for $\mu_{\rm l}=0$ with the Gaussian initial condition normalized at $1$ year, as discussed in the main text.
@@
Figure 11: Surface densities plotted against radius at different times for the magnetized $\alpha$-viscosity (solid) and the linear viscosity (dashed), similar to Fig. 4. The first case (top) assumes a Gaussian initial condition, with the normalization factor of the linear viscosity parametrization, $\nu_{0,\rm{l}}$, chosen to match the magnetized viscosity at the start time. The second case (bottom) includes a source term in the diffusion equation (Eq. 1) instead of a Gaussian initial condition.
@@
Figure 12: Light curves for near-UV, far-UV and X-ray bands (left to right) with the magnetized $\alpha$-viscosity (red) and the linear viscosity (dashed black, blue, brown), similar to Fig. 5. The first case (top) assumes a Gaussian initial condition, with the normalization factor of the linear viscosity parametrization, $\nu_{0,\rm{l}}$, chosen to match the magnetized viscosity at the initial condition. The second case (bottom) includes a source term in the diffusion equation (Eq. 1) instead of a Gaussian initial condition.
@@
Figure 13: The effective temperature $T_{\rm eff}$ vs radius at different times for the magnetized viscosity (solid) and the linear viscosity (dashed). The first case (top) assumes a Gaussian initial condition, where the normalization factor for the linear parametrization $\nu_{0,\rm l}$ is determined by matching the disk’s outer radius after 1 year. The second case (middle) also assumes a Gaussian initial condition, but the normalization factor is chosen to match the magnetized viscosity at the initial condition. The third case (bottom) includes a source term in the diffusion equation (Eq. 1) instead of using a Gaussian initial condition.

### v2

@@ 5 Conclusions
Standard accretion disk models are viscously and thermally unstable in the inner regions of the disk, where radiation pressure dominates over gas pressure (Pringle+1973; Shakura&Sunyaev1976; Lightman1974; Piran1978). However, these instabilities are not so far observed in TDEs. A common theoretical approach to stabilize the disk is to modify the viscosity parametrization (Sakimoto&Coroniti1981; Taam&Lin1984; Mummery2019), though the most widely used parametrizations are often ad hoc.
In this paper, we examine the $\alpha$-disk model (Shakura&Sunyaev1973), under the assumption that magnetic fields are the dominant source of pressure. Given the presence of MRI, we expect the disk to be highly magnetized. Previous research has shown that magnetic fields can indeed stabilize the disk against radiation pressure instabilities (Begelman&Pringle2007; Sadowski2016; Jiang+2019; Huang+2023; Kaur+2023).
@@
We examined various values for the power law of the linear viscosity and found that, both the surface density at late times and the light curves may look very similar when the power law of the linear viscosity is chosen carefully. Specifically, the surface density at late times matches well when the linear viscosity follows a power law with $\mu_{\rm l}=5/9$. Additionally, we found that the time evolution of the UV luminosity for the magnetized viscosity scales as $L_{\rm m,UV}\propto t^{-5/6}$, and that the light curves for the linear viscosity fortuitously follow the same time evolution when the power law is set to $\mu_{\rm l}=0$.
We fit our model to the well-observed optical and UV light curves of the TDE ASASSN-14li. The evolution of its light curves indicates that the viscous spreading in ASASSN-14li is slow, leading previous studies to propose mechanisms such as angular momentum loss in a magnetized wind (Wen+2023) to limit the disk’s outer spreading. Our results show that the best-fit Shakura-Sunyaev $\alpha$ parameter falls within a theoretically reasonable range, requiring no additional physical effects.
@@
In this paper, we studied accretion disks in a 1D framework, assuming a single pressure component due to magnetic fields. In reality, radiation pressure may also be non-negligible (Jiang+2019) and should be considered alongside magnetic pressure for a more comprehensive model. Furthermore, while we used an approximate estimation for the magnetic field strengths based on an MRI saturation criterion (Pessah&Psaltis2005; Begelman&Pringle2007) which is consistent with recent radiation MHD simulations (Jiang+2019; Mishra+2022; Huang+2023), alternative saturation criteria (Oda+2009; Begelman&Armitage2023) and other MHD instabilities (Das+2018) deserve further investigation and could potentially influence our results. Nevertheless, the magnetized viscosity parametrization adopted in this work remains both simple and more physically motivated compared to other commonly used parametrizations.
@@ Appendix A Separability of Disk Equation
The evolution equation (Eq. 1) with a magnetized $\alpha$-viscosity (Eq. 5) is a nonlinear partial differential equation. Here we show that Eq.˜1 can be reduced to two ordinary differential equations: (i) a trivial time-dependent equation and (ii) a nonlinear Emden-Fowler equation. To the best of our knowledge, this has not been demonstrated previously despite the wide applicability of Eq. 1 in high energy astrophysics. There are, however, more specific investigations of separability of Eq. 1 in certain limits, such as self-similar solutions (Rafikov2016).
@@
where for the magnetized $\alpha$-viscosity $m=2/7$ and $n=5/7$. We use the following transformation from Pringle1991: $x\equiv R^{1/2}$, $\tau\equiv 3\nu_{0}t/4$ and $S(\tau,x)\equiv\Sigma R^{3/2}$. Eq.˜1 then becomes:
@@
The time-dependent equation is trivial and can be solved easily. For $m=0$, the solution to Eq.˜1 has been determined previously (e.g. Metzger+2008), so we focus here on the case $m\neq 0$. In this case, the solution is $T(\tau)=(-m\lambda\tau+C)^{-1/m}$ where $C$ is an arbitrary constant of integration.
The spatial equation is a nonlinear ordinary differential equation, which through a change of variables $w(x)\equiv X^{m+1}x^{2n-3m-2}$ transforms into an Emden-Fowler [^10] The Lane-Emden equation is a well-studied special case of the Emden-Fowler equation. equation (zaitsev&Polyanin2002):
@@
For the magnetized $\alpha$-viscosity we employ, the power of $w$ is $7/9$, and the power of $x$ is $10/9$. Although exact solutions exist for specific power-law values (zaitsev&Polyanin2002), no exact solution has yet been found for these power-law indices.
We have explored other astrophysically interesting power-law viscosity parametrizations. We calculated the values for every combination of gas, radiation, and magnetic pressures ($P\in\left\{P_{\rm gas},P_{\rm rad},P_{\rm mag}\right\}$), with every combination of electron scattering, Kramers, and dust opacities ($\kappa\in\left\{\kappa_{\rm es},\kappa_{\rm Kr},\kappa_{\rm dust}\right\}$). The power-law values of the exact solutions (both continuous and discrete), as well as those of astrophysical interest, are presented in Fig.˜9. Unfortunately, no exact solutions currently exist for the values relevant to our study. If in the future such solutions are found, they could provide greater physical intuition for the time evolution of viscously spreading accretion disks.
@@ Appendix B Fitting the Self-Similar Solution
Using the self-similar solution is simpler and more efficient than numerically integrating the disk equations. In §2, we demonstrated that the self-similar solution closely approximates the numerical solution at late times and large radii. In this appendix, we provide a fitting function for the free parameter $\nu_{0,\rm{S}}$ in Eq.˜12. For each value of the $\alpha$ parameter, SMBH mass $M_{\bullet}$, and stellar mass $M_{\star}$ satisfying $M_{\bullet}<M_{\rm Hills}$, the self-similar free parameter can be determined using:
@@
Fig.˜10 shows the relative error in the light curves at a UV wavelength of $\lambda=250\rm{nm}$, between the numerical and the self-similar solutions:
$\rm{RelativeError(t,\alpha,M_{\bullet},M_{\star})}=100\frac{L_{{\color[rgb]{1,0,0}\textbf{}}f,\rm{num}}-L_{{\color[rgb]{1,0,0}\textbf{}}f,\rm{S}}}{L_{{\color[rgb]{1,0,0}\textbf{}}f,\rm{num}}}\,[\%]$ (B2)
where $t$ is the age of the TDE, $L_{{\color[rgb]{1,0,0}\textbf{}}f,\rm{num}}$ is the spectral luminosity calculated from the numerical solution, and $L_{{\color[rgb]{1,0,0}\textbf{}}f,\rm{S}}$ is the spectral luminosity from the self-similar solution. The relative error is larger for SMBH masses near the Hills mass but remains below an order of magnitude in other regions.
@@ Appendix C Comparison To Linear Viscosity with Additional Models
In this appendix, we present additional results. We compare the magnetized $\alpha$-disk with the linear viscosity model, as was done in §3, for two additional scenarios. In the first scenario, we retain the Gaussian initial condition but normalize the linear viscosity prefactor, $\nu_{0,\rm l}$, in the initial conditions, rather than matching it to the magnetized disk after 1 year. Specifically, we set: $\nu_{0,\rm l}=\nu_{\rm mag}\left(r_{\rm c},\Sigma_{0}\right)/r_{\rm c}^{\mu_{\rm l}}$ where $\Sigma_{0}=\Sigma\left(R=r_{\rm c},t=0\right)$ is the initial surface density at the peak of the Gaussian. We compute results for various power-law values $\mu_{\rm l}=\left\{1,5/9,0,-3/2\right\}$.
@@
In the second scenario, we add a source function to Eq.˜1 instead of using a Gaussian initial condition (while retaining the original normalization $\nu_{0,\rm{l}}$). This approach reflects the slow accumulation of mass from the bound debris into the accretion disk, at a rate given by Eq. 9. For this case, we compute results for $\mu_{\rm l}=\left\{3/2,5/9,0,-3/2\right\}$.
@@
In Figs.˜11, 12, 13 and 14, we explore these non-fiducial models by plotting the surface density and effective temperature as functions of disk radius, as well as the light curves. For all figures, we use same SMBH, star, and disk parameters as in §3. The best match between the magnetized and linear viscosity models is achieved for $\mu_{\rm l}=0$ with the Gaussian initial condition normalized at $1$ year, as discussed in the main text.
@@
Figure 11: Surface densities plotted against radius at different times for the magnetized $\alpha$-viscosity (solid) and the linear viscosity (dashed), similar to Fig.˜4. The first case (top) assumes a Gaussian initial condition, with the normalization factor of the linear viscosity parametrization, $\nu_{0,\rm{l}}$, chosen to match the magnetized viscosity at the start time. The second case (bottom) includes a source term in the diffusion equation (Eq. 1) instead of a Gaussian initial condition.
@@
Figure 12: Light curves for near-UV, far-UV and X-ray bands (left to right) with the magnetized $\alpha$-viscosity (red) and the linear viscosity (dashed black, blue, brown), similar to Fig.˜5. The first case (top) assumes a Gaussian initial condition, with the normalization factor of the linear viscosity parametrization, $\nu_{0,\rm{l}}$, chosen to match the magnetized viscosity at the initial condition. The second case (bottom) includes a source term in the diffusion equation (Eq. 1) instead of a Gaussian initial condition.
@@
Figure 13: The effective temperature $T_{\rm eff}$ vs radius at different times for the magnetized viscosity (solid) and the linear viscosity (dashed). The first case (left) assumes a Gaussian initial condition, where the normalization factor for the linear parametrization $\nu_{0,\rm l}$ is determined by matching the disk’s outer radius after 1 year. The second case (right) also assumes a Gaussian initial condition, but the normalization factor is chosen to match the magnetized viscosity at the initial condition.
Figure 14: Same as Fig.˜13, but includes a source term in the diffusion equation (Eq. 1) instead of using a Gaussian initial condition.